# No.2 Linggong Road, Dalian, Liaoning, 116024, P.R.China

### v1

@@ 4.5 Generalization across feature sets
To evaluate the generalization ability, we train the LCF models on the Mock5 and Mock6 datasets. Mock5 consists of a typical combination of features, and Mock6 introduces random reordering of features to test whether the model remains robust to different feature arrangements. The learned cuts for Mock5 are shown in figure 22 (parallel LCF) and figure 23 (sequential LCF) with corresponding learned importance scores shown in figure 25 and figure 25. The figures related to Mock6 are 22, 23, 25, and 25..
Figure 22: Learned cuts from the parallel LCF on the Mock5 dataset.
Figure 23: Learned cuts from the sequential LCF on the Mock5 dataset.
Figure 24: Learned importance from the parallel LCF on the Mock5 dataset.
Figure 25: Learned importance from the sequential LCF on the Mock5 dataset.
Figure 26: Learned cuts from the parallel LCF on the Mock6 dataset.
Figure 27: Learned cuts from the sequential LCF on the Mock6 dataset.
Figure 28: Learned importance from the parallel LCF on the Mock6 dataset.
Figure 29: Learned importance from the sequential LCF on the Mock6 dataset.
@@
In the parallel strategy for Mock5, the learned cuts are consistent with those observed in previous Mock datasets except for feature $x_{7}$ whose cut covers the full range due to its lack of signal-background separation as shown in figure 22. Unlike the seemingly random cuts for $x_{7}$ in section 4.3, this result more clearly demonstrates that the model directly ignores such a feature. This matches its importance score which is zero in figure 25. The relative importance among $x_{1}$ to $x_{4}$ remains similar to Mock1. Feature $x_{9}$ is more important than $x_{5}$, which aligns with their signal-background separation. In the sequential strategy, it clearly shows that the cuts for $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ effectively separate signal from background. By contrary, $x_{5}$’s cut has limited separation and $x_{7}$ and $x_{9}$ cover the full range, which matches their zero importance scores in figure 25.
@@
The results from the reordered dataset Mock6 are interesting. In the parallel strategy, since every feature is treated separately, the importance scores are nearly identical to those in Mock5 (figure 29). By contrast, in the sequential strategy, the presence of a less informative feature like $x_{7}$ early in the sequence doesn’t prevent the model from converging correctly. The cut boundaries for $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ remain clear and effective while $x_{7}$ and $x_{9}$ cover the full range as in Mock5. The cut for $x_{5}$ appears less constrained than in Mock5, leading to a small non-zero importance in figure 29, though its separation remains limited.
@@ 4.6 Benchmark on diboson classification
Finally, we evaluate the LCF model on the real diboson dataset in both strategies and also compare its performance with other baseline models. The learned cuts are shown in figure 30 (parallel LCF) and figure 31 (sequential LCF). The corresponding learned importance scores are shown in figure 33 and figure 33. The performance is summarized in table 2. The table presents true positive (TP, the number of retained signal events), false positive (FP, the number of inevitable background events), accuracy (the number of correctly tagged events), precision (the ratio of correctly tagged signal), and significance which is calculated by $\text{TP}/\sqrt{\text{FP}}$.
@@
Figure 32: Learned importance from the parallel LCF on the diboson dataset.
Figure 33: Learned importance from the sequential LCF on the diboson dataset.
@@
In the parallel strategy, the learned cuts are intuitively placed at the boundaries where signal and background distributions diverge, effectively separating the two classes across all observables. In contrast, in the sequential strategy, the first two cuts on $M_{\mathrm{jet}}$ and $C_{2}^{\beta=1}$ filter out most background events. This causes the signal and background to almost completely overlap in the distributions of the later observables ($\tau_{21}^{\beta=1}$, $C_{2}^{\beta=2}$, $D_{2}^{\beta=1}$, $D_{2}^{\beta=2}$), making their cuts less effective. To confirm whether these overlapped observables are truly negligible, we examine the learned importance scores. In both strategies, all observables are assigned fair importance scores, with none close to zero, indicating that there are still signal samples in the long-tail regions of these distributions. Comparing the scores, $D_{2}^{\beta=2}$ ranks second in the parallel strategy and is more important than $D_{2}^{\beta=1}$, which aligns with the smaller overlap area in its distribution. In the sequential strategy, this trend persists, though the importance of later observables is generally reduced. Notably, $\tau_{21}^{\beta=1}$ is less important than both ${C_{2}^{\beta=1}}^{\beta=1}$ and $C_{2}^{\beta=2}$ in both strategies. This is likely because importance is more related to the absolute area of separation as observed in the figures.
Among all the tested models, the LCF model, as a cut-based method, exhibits the highest precision despite a lower accuracy. The parallel strategy effectively minimizes background events to a small fraction, albeit at the expense of a greater signal loss compared to other models. In contrast, the sequential strategy behave more like a classifier, enhancing accuracy by systematically checking each distribution while considering the effects of prior cuts.
We further investigate performance differences through two additional experiments. The first experiment only keeps the observables with the initial importance scores above the average (1/7) namely $M_{\text{jet}}$, $D_{2}^{\beta=1}$, and $D_{2}^{\beta=2}$. The second experiment excludes highly-correlated observables, retaining $M_{\text{jet}}$, $C_{2}^{\beta=2}$, $D_{2}^{\beta=2}$, and $\tau_{21}^{\beta=1}$, as informed by the correlation figure. The results of these experiments are shown in table 3 and table 4.
Model TP FP Accuracy Precision Significance
BDT 42046 7422 0.8567 0.8500 488.0499
MLP 42408 7532 0.8597 0.8492 488.6441
LCF (Parallel) 34266 4125 0.8029 0.8926 533.5209
LCF (Sequential) 40027 7078 0.8366 0.8497 475.7708
Table 3: Performance metrics of different models on the first dataset.
Model TP FP Accuracy Precision Significance
BDT 41066 6874 0.8655 0.8566 495.3106
MLP 40915 6568 0.8675 0.8617 504.8542
LCF (Parallel) 28111 3751 0.7376 0.8823 458.9895
LCF (Sequential) 40611 7304 0.8540 0.8476 475.1856
Table 4: Performance metrics of different models on the second dataset.
Retaining observables with higher-than-average importance leads to clear improvements: the parallel LCF model achieves a significance of 533 in table 3 (up from 453.95 in table 2), surpassing the 488 of both the BDT and MLP. In contrast, removing highly-correlated observables has little effect. This prove once again that the learnable importance is not only a metrics but also a practical standard to select observables.
@@ 5 Conclusion
In this work, we introduce the Learnable Cut Flow (LCF), a novel neural network framework that bridge the interpretability of traditional cut-based methods with the adaptability of modern machine learning techniques in high-energy physics. By transforming the manual process of cut selection into a fully differentiable, data-driven operation, LCF offers a transparent and actionable alternative to opaque black-box models. Two widely-used strategies are also implemented via mask operations making LCF a fully interpretable model. We also propose the Learnable Importance to dynamically adjust each observable’s contribution to the final classification results.
Testing on six synthetic mock datasets and a realistic diboson vs. QCD dataset demonstrated LCF’s versatility and effectiveness. The model accurately identifies optimal cut boundaries across a range of feature distributions, including left, right, middle, and edge cases, while dynamically adapting to varying degrees of signal-background separation. The Learnable Importance mechanism successfully assigns higher weights to features with stronger discriminative power, effectively suppresses redundant or highly correlated observables, and remains robust to feature ordering. On the diboson dataset, LCF initially lagged behind boosted decision trees (BDT) and multilayer perceptrons (MLP) in raw performance. However, by pruning less critical observables guided by learned importance, the parallel LCF achieved a significance of 533.52, surpassing the baselines (488.05 for BDT and 488.64 for MLP), underscoring the practical utility of this metric.
LCF’s ability to emulate human cut-searching strategies while embedding them within a trainable framework marks a significant step forward. It retains the simplicity and physicist-friendly insights of traditional cut flows, yet harnesses the power of neural networks to handle complex, high-dimensional data. This hybrid approach not only enhances classification performance but also provides interpretable outputs—cut boundaries and importance scores—that can guide observable selection in real-world analyses.

### v3

@@ 4.5 Generalization across feature sets
To evaluate the generalization ability, we train the LCF models on the Mock5 and Mock6 datasets. Mock5 consists of a typical combination of features, and Mock6 introduces random reordering of features to test whether the model remains robust to different feature arrangements (here the order of features is $x_{1},x_{2},x_{7},x_{3},x_{5},x_{4},x_{9}$). The learned cuts on Mock5 are shown in figure 18 (parallel LCF) and figure 19 (sequential LCF) with corresponding learned importance scores shown in figure 20. The figures related to Mock6 are 2122, and 23.
@@
In the parallel strategy for Mock5, the learned cuts are consistent with those observed in previous Mock datasets except for feature $x_{7}$. The difference here is that the weighted baseline defined in eq. (34) is lower than in the Mock3 dataset, which includes another redundant feature $x_{8}$. So the importance of $x_{7}$ is compressed much harder leading to zero. Unlike the randomly fixed cuts on $x_{7}$ in section 4.3, the learned cuts on $x_{7}$ here are totally random due to the zero importance score which leads to no adjustment to the gradient and ends up with a quite loose cut. The relative importance among $x_{1}$ to $x_{4}$ remains similar to Mock1. Feature $x_{9}$ is more important than $x_{5}$, which aligns with their signal-background separation. In the sequential strategy, it clearly shows that the cuts on $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ effectively separate signal from background. In contrast, the cuts on $x_{5}$, $x_{7}$ and $x_{9}$ cover the full range, which matches their zero importance scores in figure 20. The first four cuts are effective and dominate the subsequent cuts, which is consistent with the results in Mock1. The weak feature $x_{5}$ and highly-correlated feature $x_{9}$ don’t provide any information to the final classification according to their large boundary.
Figure 18: Learned cuts from the parallel LCF on the Mock5 dataset. The error bars are not shown due to the small error or no meaning under the zero importance.
Figure 19: Learned cuts from the sequential LCF on the Mock5 dataset. The error bars are not shown due to the small error or no meaning under the zero importance.
Figure 20: Learned importance from the parallel (left) and sequential (right) LCFs on the Mock5 dataset. The baseline indicates the minimum importance (default: 5% of average importance = $1/F\times 0.05$), below which the features are ignored during inference.
Figure 21: Learned cuts from the parallel LCF on the Mock6 dataset. The error bars are not shown due to the small error or no meaning under the zero importance.
Figure 22: Learned cuts from the sequential LCF on the Mock6 dataset. The error bars are not shown due to the small error or no meaning under the zero importance.
Figure 23: Learned importance from the parallel (left) and sequential (right) LCFs on the Mock6 dataset. The baseline indicates the minimum importance (default: 5% of average importance = $1/F\times 0.05$), below which the features are ignored during inference.
@@
The results from the reordered dataset Mock6 are interesting. In the parallel strategy, since every feature is treated separately, the importance scores are nearly identical to those in Mock5 (figure 23). In contrast, in the sequential strategy, the presence of a less informative feature like $x_{7}$ early in the sequence doesn’t prevent the model from converging correctly. The cut boundaries for $x_{1}$, $x_{2}$, $x_{3}$, $x_{4}$ remain clear and effective while $x_{7}$ and $x_{9}$ cover the full range as in Mock5. The cut on $x_{5}$ is the only one changed compared to Mock5. The key difference is $x_{4}$ is applied later than $x_{5}$ and $x_{5}$ doesn’t generate that large loss like $x_{7}$, so the importance of $x_{5}$ is not compressed that hard like in Mock5. The change of applying order and the nature of the feature itself are the main reason for the difference.
Model TP FP Accuracy Precision Significance
LCF (Parallel) $16532\pm 51$ $\pm 0 66$$6.7\pm 0.1\%$ $100.0\pm 0.0\%$ $52.1\pm 1.1$
LCF (Sequential) $39769\pm 40$ $1571\pm 25$ $88.3\pm 0.0\%$ $96.2\pm 0.1\%$ $7.8\pm 0.1$
Table 6: Performance metrics of LCF models on the Mock5 dataset.
Model TP FP Accuracy Precision Significance
LCF (Parallel) $16245\pm 94$ $\pm 0 66$$6.1\pm 0.1\%$ $100.0\pm 0.0\%$ $50.8\pm 1.0$
LCF (Sequential) $40116\pm 22$ $1552\pm 13$ $88.4\pm 0.0\%$ $96.3\pm 0.0\%$ $7.9\pm 0.0$
Table 7: Performance metrics of LCF models on the Mock6 dataset.
Tables 6 and 7 show clearly that the results of the sequential strategy are more stable than those of the parallel strategy. Specifically, the sequential strategy demonstrates lower variance across multiple training runs, as evidenced by smaller standard deviations in the significance values: $7.8\pm 0.1$ and $7.9\pm 0.0$ for Mock5 and Mock6 respectively, compared to $52.1\pm 1.1$ and $50.8\pm 1.0$ for the parallel strategy. Meanwhile, the parallel strategy achieves far better significance than the sequential strategy following the same logic stated in section 4.4.
@@ 4.6 Benchmark on diboson classification
Finally, we evaluate the LCF model on the real diboson dataset in both strategies and also compare its performance with other baseline models. The learned cuts are shown in figure 24 (parallel LCF) and figure 25 (sequential LCF). The corresponding learned importance scores are shown in figure 26. The performance is summarized in table 8.
Figure 26: Learned importance from the parallel (left) and sequential (right) LCFs on the diboson dataset. The baseline indicates the minimum importance (default: 5% of average importance = $1/F\times 0.05$), below which the features are ignored during inference.
@@
In the parallel strategy, the learned cuts are intuitively placed at the boundaries where signal and background distributions diverge, effectively separating the two classes across all observables. In contrast, in the sequential strategy, the first two cuts on $M_{\mathrm{jet}}$ and $C_{2}^{\beta=1}$ filter out most background events. This causes the signal and background to almost completely overlap in the distributions of the later observables ($C_{2}^{\beta=2}$, $D_{2}^{\beta=1}$, $D_{2}^{\beta=2}$, $\tau_{21}^{\beta=1}$), making their cuts less effective. To confirm whether these overlapped observables are truly negligible, we examine the learned importance scores. In both strategies, all observables are assigned fair importance scores, with none close to zero. Comparing the scores, $D_{2}^{\beta=2}$ ranks second in the parallel strategy and is more important than $D_{2}^{\beta=1}$, which aligns with the smaller overlap area in its distribution. In the sequential strategy, this trend persists, though the importance of later observables is generally reduced. Notably, $\tau_{21}^{\beta=1}$ is equally important as $C_{2}^{\beta=2}$ and non-zero in both strategies even if its cut covers the full range in the sequential strategy. According to eq. (34) and the results from Mock3 (redundant features) and Mock4 (high correlated features), importance scores tend toward zero when features generate relatively high loss compared to others (like $x_{7}$ and $x_{8}$) or when highly correlated features have decreasing importance based on their application order. $\tau_{21}^{\beta=1}$ represents a different case: its signal-background separation is at a similar level to other observables, and during sequential LCF training, it generates moderate loss compared to truly redundant features. Additionally, since it is not highly correlated with other observables, its importance score does not decrease like $D_{2}^{\beta=1}$ and $D_{2}^{\beta=2}$.
The training time of the LCF model is at the same level as the BDT and MLP models, which is lightweight to train compared to modern deep neural networks. The significance of the LCF models, however, is slightly worse than the BDT and MLP models because they are simpler and shallower networks essentially. The parallel strategy this time doesn’t outperform the sequential strategy in terms of significance. The parallel strategy’s more conservative cut boundaries, while effective in the Mock datasets, here incur signal efficiency costs that outweigh the background reduction benefits in this diboson dataset. This demonstrates the fundamental trade-off inherent in the parallel strategy’s approach. From this perspective, the sequential strategy is always more stable (lower variance of metrics across training runs) and more accurate (as a classifier with higher accuracy). However, the parallel strategy gains advantages due to the cuts that achieve lower classification accuracy—it always filters out more events. Combined with results from Mock datasets, such inaccuracy doesn’t lead to a consistent significance improvement. The sequential strategy is a better choice as a classifier, just like all other neural networks that are trained to reduce the loss designed for classification. On the other side, the parallel strategy tends to achieve better significance values in most cases, though the optimal choice depends on the specific dataset characteristics. Both strategies can be quickly validated in LCF and the better one can be chosen.
@@ 5 Conclusion
In this work, we introduce the Learnable Cut Flow (LCF), a novel neural network framework that bridges the interpretability of traditional cut-based methods with the adaptability of modern machine learning techniques in high-energy physics. By transforming the manual process of cut selection into a fully differentiable, data-driven operation, LCF offers a transparent and actionable alternative to opaque black-box models. Two widely-used strategies are also implemented via mask operations making LCF a fully interpretable model. We also propose the Learnable Importance to dynamically adjust each observable’s contribution to the final classification results.
Testing on six synthetic mock datasets and a realistic diboson vs. QCD dataset demonstrated LCF’s versatility and effectiveness. The model accurately identifies optimal cut boundaries across a range of feature distributions, including left, right, middle, and edge cases, while dynamically adapting to varying degrees of signal-background separation. The Learnable Importance mechanism successfully assigns higher weights to features with stronger discriminative power, effectively suppresses redundant or highly correlated observables, and remains robust to feature ordering. Although the LCF is less performant than the baseline models on the diboson dataset, it is as lightweight to train as they are, requiring fewer computational resources than modern deep neural networks. This enables the full set of optimized cuts to be obtained with reduced analyst effort and at minimal cost.
LCF’s ability to emulate human cut-searching strategies while embedding them within a trainable framework marks a significant step forward. It retains the simplicity and physicist-friendly insights of traditional cut flows, yet harnesses the power of neural networks to handle complex, high-dimensional data. This hybrid approach not only enhances classification performance but also provides interpretable outputs—cut boundaries and importance scores—that can guide observable selection in real-world analyses. Future extensions of LCF could include optimization directly against signal significance—particularly valuable for BSM searches with low cross sections—and a mechanism to learn optimal feature ordering, enabling full automation while preserving interpretability.